### v1

@@ 4.1 Area Attributes Fusion Module
Similar areas usually have similar traffic characteristics, e.g., residential locations usually have relatively high demand for charging at night. The type and number of POIs can be a good indicator of the area’s utilization attributes, but it is not in the same form as demand data and cannot be used directly as an input to the model. We use POIs as the basis for area clustering to learn the similar pattern in areas groups. To enhance the accuracy of clustering, Term Frequency-Inverse Document Frequency (TF-IDF) and K-means are employed. TF-IDF is a common weighting technique used in information retrieval and mining to evaluate the importance of a word to a document. We treat a POI category as a word and an area as a document, and use this method to assess the importance of each category for the areas.
Term Frequency (TF) indicates the normalized frequency of occurrence of a term in the text, as shown in Eq. (2).
${\rm{TF}}_{i,j}=\frac{f_{i,j}}{\sum_{k}f_{k,j}}$ ((2))
where $\rm{TF}_{i,j}$ denotes the term frequency of word $i$ in document $j$, and $f_{i,j}$ denotes the number of times word $i$ appears in document $j$.
Note that some generic words do not contribute much to the theme, but rather less frequent words that convey the theme of the text, so it is not appropriate to use TF simply. The same is true for POI fusion, where some of the high-frequency POI types may not be fully representative of the area’s usage characteristics. The Inverse Document Frequency (IDF) indicates the prevalence of the word. The fewer the documents containing the word, the larger the IDF, indicating that the word has good category differentiation. The calculation of IDF is shown in Eq. (3).
${\rm IDF}_{i,j}=\log{\frac{N_{d}}{1+\left|j:i\in j\right|}}$ ((3))
where $\rm{IDF}_{i,j}$ denotes the inverse document frequency of word $i$ in document $j$, $N_{d}$ denotes the number of documents, and $\left|j:i\in j\right|$ denotes the number of documents containing word $i$.
Taken together, a high TF-IDF is expected if the frequency of a POI type is high in a specific area and the frequency of that POI is low in the entire city area. The calculation of TF-IDF is shown in Eq. (4).
${\rm{TF\text{-}IDF}}_{i,j}={\rm TF}_{i,j}\times{\rm IDF}_{i,j}$ ((4))
where $\rm{TF\text{-}IDF}_{i,j}$ denotes the Term Frequency-Inverse Document Frequency of word $i$ in document $j$.
From the above we obtain the sequence $\rm{TF\text{-}IDF}_{:,j}$ of the importance of each POI type for each area. Then, we employ K-means clustering [50] to divide the areas into a specific number of groups (denoted as $C$) based on TF-IDF.
In order to fuse features within groups, hypergraph attention mechanism is introduced. We consider all areas of the entire city as a hypergraph with incidence matrix $\bm{M}\in\left\{0,1\right\}^{\left|\nu\right|\times\left|\varepsilon\right|}$, where $\nu$ denotes the vertex set and $\varepsilon$ denotes the hyperedge set. Besides, we consider each area group as a hyperedge $\bm{e}_{i}$, which belongs to $\varepsilon$. Hyperedge $\bm{e}_{i}$ connects specific areas ${\left\{\bm{n}_{j}\mid n_{j}\in e_{i}\right\}}$. Attentive hypergraph neural network is an improvement of hypergraph convolution network, which is divided into two steps.
@@
To avoid repetition, we take the first step as an example, showing how the update embedding $\bm{e}^{(1)}_{i}$ of a hyperedge is aggregated from itself and its connected nodes. The computational process is essentially the same as the graph attention mechanism, where features are concatenated and then feature-aware soft attention mechanism is used to determine the importance between nodes. The specific calculation process is shown in Eq. (6).
@@ 4.3 Temporal Feature Fusion Module
To accurately predict the EV charging demand, it is worthwhile to take into account external information (e.g., price, temperature) that can influence the user’s choice of charging service. Considering auxiliary information in the model is expected to effectively capture dynamic changes and reduce lags in prediction. The data structure of some external information is the same as the time-series demand data. How to incorporate this information into the model and model the temporal effects to improve prediction accuracy remains a challenge.
@@
Fig. 3 illustrates the structure of three building blocks of the proposed model. First, we introduce Gated Residual Network (GRN), as shown in Fig. 3(a), in order to allow the model to flexibly handle nonlinear relationships when needed. The specific calculation process of GRN is shown in Eq. (11).
@@
Inspired by soft attention mechanism, we improve a variable selection network as shown in Fig. 3(b). Auxiliary information and historical demand data used for urban EV charging demand prediction are inputted. The specific contribution of a variable is determined by the network, i.e., variable selection network enables to emphasize variables that are critical to the prediction problem and ignore variables that are not. The input data $\bm{\xi}$ (including high-dimensional features $\bm{H}$, and other dynamics $\bm{U}_{s}$) are processed independently by two GRNs. A softmax function after one of the GRNs determines the variable selection weights, which are then multiplied with the embedded features. The calculation process of the variable selection network is shown in Eq. (13).
@@
where $\bm{\xi}_{1}$ is the embedded features after the first step calculation, $\bm{v}_{1}$ is the variable selection weights, and $\bm{\xi}_{2}$ denotes the output of variable selection network, which combines multiple features.
Transformer utilizes self-attention mechanism to model sequences, breaking through the limitation that recurrent neural networks cannot be computed in parallel, and becoming a research hotspot for time series tasks [12]. We adopt the framework of Transformer encoder and improve it by utilizing GRN and Gated Attention Network. Similarly to self-attention, we process the output of variable selection network (i.e., $\bm{\xi}_{2}$) independently with three GRNs to obtain the query, key, and value matrix $\bm{Q}$, $\bm{K}$, and $\bm{V}$.
The attention mechanism assigns weights to each input, and even irrelevant units are assigned small weights. This leads to two problems: 1) relevant units are assigned insufficient attention weights, and 2) units without contributions are assigned weights, leading to performance degradation [34]. Considering that in long sequences not every interval is sufficiently helpful for prediction, we develop gated temporal attention, in which we perform a Scaled Dot-Product Attention operation, but Gumbel-Softmax [53] is used as weight activation function instead of Softmax, as shown in Eq. (14) and Fig. 3(c).
@@ 5 Experiments
In this section, we first present the data description and the specific setup of the experiment. Then, we compare the proposed model with current state-of-the-art models to empirically illustrate the performance of AHMDF. The model generalizability is also illustrated. Furthermore,we conduct ablation experiments to show that each module and each factor can provide a contribution to improving the prediction accuracy. Finally, the impact of hyperparameters is analyzed experimentally.
@@ 5.1 Data Description
The proposed model AHMDF is validated on a benchmark dataset collected from June 19 to July 18, 2022 in Shenzhen, China. We divide the city into areas and update the occupancy rate of the counted charging piles within the area every 5 minutes. Consequently we can obtain 8640 timestamps for the month being counted. The areas division and the number of charging piles included in each area are shown in this link [^1] https://github.com/IntelligentSystemsLab/ST-EVCDP. We divide the training set, validation set and test set by 6:1:3 in chronological order, i.e., 19 June to 6 July for training, 7 July to 9 July for validation and 10 July to 18 July for test. A total of 450,175 POI data are collected from Amap [^2] https://lbs.amap.com/. POIs include a total of 14 categories such as restaurant, shopping, scenic areas, etc. Temperature data is also collected from a weather station in Shenzhen, which has a sampling interval of 30 minutes. We use linear interpolation to enable a 5-minute interval in the temperature data, making it the same as the demand data.
@@ 5.2 Experimental Setup
The hyperparameters are chosen after a detailed search experiment and we provide hyperparameter analysis in Subsection 5.6. The final experimental setup we chose is as follows. First, the lookback window size $\tau$ is set to 12, i.e. we provide the model with 60 minutes of historical information. Second, in K-means clustering, we cluster the area into 10 classes. Third, the temperature coefficient $\varpi$ is set to 1.5 to not only provide sparse attention to temporal information, but also to avoid focusing attention on one-hot vector. Fourth, the number of encoder components $N_{E}$ is set to 2. Fifth, the maximum epoch numbers of training process is set to 2000, and model training will early stop if the loss of the validation set does not decrease within 50 epochs. Sixth, Mean Square Error (MSE) is used as the loss function for all comparison models. Seventh, we use Adam [54] as an optimizer with a mini-batch size of 512. Seventh, we employ four metrics, i.e., Root Mean Squared Error (RMSE), R-Square Coefficient ($\rm{R}^{2}$), Relative Absolute Error (RAE), and Mean Absolute Error (MAE) to illustrate the performance of our proposed model. Finally, all experiments in this study are accomplished on a Windows personal computer containing a NVIDIA RTX 4000 GPU. We share the code for our proposed model AHMDF in this link. [^3] https://github.com/kuanghx3/AHMDF.
@@ 5.3 Comparison Experiment
In this subsection, the proposed model AHMDF will be compared with representative models to illustrate the state-of-the-art performance. The models for comparison includes 3 sequential neural networks, 2 spatial neural networks and 5 spatio-temporal neural networks, which are described in detail as follows:
@@
  • TGCN [30]: A neural network-based traffic prediction method combining graph convolutional networks (GCN) and gated recurrent units (GRU).
@@
  • BEGAN [33]: An air traffic demand prediction network combining long short-term memory (LSTM) and graph attention mechanisms (GAT).
  • HSTGCN [18]: A heterogeneous spatio-temporal graph convolutional network combining GCN and GRU for predicting EV charging demand at different spatio-temporal resolutions. In particular, it considers POI classification in the model.
  • PIAST [21]: A spatio-temporal electric vehicle charging demand prediction model using convolutional neural network and physics-informed neural network to fuse price information, which is our previous work.
@@
In Table 1, we demonstrate the performance of the above comparison models and AHMDF in four prediction intervals (i.e., 15, 30, 45, and 60 minutes). Although the model does not achieve the best for every prediction interval and every metric, the enhancement is still substantial. Compared to ten representative methods, our proposed method AHMDF improves about 9.35% in RMSE, 10.01% in RAE, and 9.89% in MAE on average, and achieves the best fit in $\rm R^{2}$. Moreover, AHMDF outperforms the ten comparison models by 16.60%, 9.09%, 7.93% and 7.94% on average in 15, 30, 45, and 60 minutes prediction tasks, respectively. AHMDF performs well on relatively long time prediction (e.g., 60min), which is a challenge in prediction problems, demonstrating the contribution of temporal modules (e.g., improved transformer encoder) and multi-source heterogeneous information to improve the accuracy of EV charging demand prediction. To visually demonstrate the enhancement of our proposed model, we show some prediction results of AHMDF and the previously proposed PIAST in an example area in Fig. 4. Differences between predictions and labels are plotted to illustrate this improvement clearly. The proposed method AHMDF has more accurate predictions than PIAST, with lower prediction lags when the data fluctuates dramatically (e.g., during the evening peak in this area). Furthermore, it can be found that AHMDF has an outstanding lead in long-term prediction task. This demonstrates that the effective introduction of external information such as the areas’ attributes and weather, as well as the improvement of the network structure can enhance the long-term prediction accuracy of EV charging demand.
Figure 4: Lineplots of the prediction results and errors for an example area in different prediction interval.
@@
In comparison models, firstly, GCN that only considers spatial adjacencies and lacks consideration of temporal factors is the worst. Second, complex models perform better than simple ones, because complex models can model nonlinear spatio-temporal relationships accurately. This is why current research favors delicate and complex networks, even though they may require more storage space and calculation. Thirdly, the consideration of influences on EV charging demand somehow brings about improvements in model performance. HGNN and HSTGCN, which incorporate area attribute information, have relatively outstanding performance in comparison models, illustrating that area attributes can help model spatial features from high dimensions and improve prediction accuracy by leveraging similar nodes. PIAST, which incorporates dynamic price changes to effectively infer user behavior, also achieves good prediction results.
@@
We conduct additional experiments to show that our model does not just fit on a specific dataset. We randomly drop 20% of the 247 areas to create different geographic contexts in the experiments. Further, to avoid randomness, the experiments are repeated five times and their results are averaged. On these new datasets, we selected one temporal model (i.e., LSTM) and three spatio-temporal models (i.e., BEGAN, HSTGCN, and PIAST) as comparative models to demonstrate the performance of our proposed model. The results and average results for the 4 prediction intervals are showed in Table 2. In the additional experiments, the proposed model AHMDF improves by 11.79% in RMSE, 10.62% in RAE and 11.27% in MAE on average compared to the comparison model, and achieves the best fit in $\rm R^{2}$.
@@
PIAST 2.90 4.35 5.35 6.21 4.70 92.36 84.11 76.87 69.11 80.61
AHMDF 2.69 4.08 5.05 5.83 4.41 92.81 84.50 77.09 69.63 81.01
Metrics RAE MAE

### v2

@@ 3.1 Area Attributes Fusion Module
Similar areas usually have similar traffic characteristics, e.g., residential locations usually have relatively high demand for charging at night. The type and number of POIs can be a good indicator of the area’s utilization attributes, but it is not in the same form as demand data and cannot be used directly as an input to the model. We view POIs in an area as a document, as shown in Eq. (1).
$<\rm{POI}^{j}_{1},\ \rm{POI}^{j}_{2},\ \rm{POI}^{j}_{3},......,\ \rm{POI}^{j}_{n}>$ ((1))
where $\rm{POI}^{a}$ stands for POI terms in area $j$ and $n$ denotes the number of the POI terms.
Then, we count the number of POIs in each category. We treat a POI category as a word in documents. Information retrieval and mining technique is employed to measure the importance of POI categories for an area. In order to focus on both high-frequency occurrences and significantly differentiated POI categories, Term Frequency-Inverse Document Frequency (TF-IDF)[46] is chosen, where a high TF-IDF value is expected if the frequency of a POI type is high in a specific area and the frequency of that POI is low in the entire city. The calculation of TF-IDF is shown in Eq. (2).
$u_{i,j}=\frac{f_{i,j}}{\sum_{k}f_{k,j}}\times\log{\frac{N_{j}}{1+\left|j:i\in j\right|}}$ ((2))
where $u_{i,j}$ denotes the Term Frequency-Inverse Document Frequency of POI category $i$ in area $j$, $f_{i,j}$ denotes the number of times POI category $i$ appears in area $j$, $N_{j}$ denotes the number of areas, and $\left|j:i\in j\right|$ denotes the number of areas containing POI category $i$.
From the above we obtain the sequence $U_{j}=\{u_{1,j},u_{2,j},...\}$ of the importance of each POI category for each area $j$. Then, we employ K-means clustering [47] to divide the areas into a specific number of groups (denoted as $C$) based on TF-IDF sequences $U$.
In order to fuse features within groups on a single network, hypergraph attention mechanism is introduced. We consider all areas of the entire city as a hypergraph with incidence matrix $\bm{M}\in\left\{0,1\right\}^{\left|\nu\right|\times\left|\varepsilon\right|}$, where $\nu$ denotes the vertex set and $\varepsilon$ denotes the hyperedge set. Besides, we consider each area group as a hyperedge $\bm{e}_{i}$, which belongs to $\varepsilon$. Hyperedge $\bm{e}_{i}$ connects specific areas ${\left\{\bm{n}_{j}\mid n_{j}\in e_{i}\right\}}$. Attentive hypergraph neural network is an improvement of hypergraph convolution network, which is divided into two steps.
@@
We take the first step as an example, showing how the update embedding $\bm{e}^{(1)}_{i}$ of a hyperedge is aggregated from itself and its connected nodes. The computational process is essentially the same as the graph attention mechanism, where features are concatenated and then feature-aware soft attention mechanism is used to determine the importance between nodes. The specific calculation process is shown in Eq. (4).
@@ 3.3 Temporal Feature Fusion Module
To accurately predict the EV charging demand, it is worthwhile to take into account external information (e.g., price, temperature) that can influence the user’s choice of charging service. Considering auxiliary information in the model is expected to effectively capture dynamic changes and reduce lags in prediction. The data structure of some external information is the same as the time-series demand data. In this subsection, we propose a temporal feature fusion scheme that is appropriate for urban EV charging demand prediction.
@@
Fig. 3 illustrates the structure of three building blocks of the proposed model. First, as shown in Fig. 3(a), we introduce Gated Residual Network (GRN) as a building block to allow the model to flexibly handle nonlinear relationships when needed. Unlike traditional networks that use simple linear layers, we choose GRNs in order to provide the network with more powerful understanding capabilities. GRNs have more sophisticated gating mechanisms and nonlinear activation functions compared to linear layers, and can adaptively choose nonlinear processing in response to fluctuations in charging demand. The specific calculation process of GRN is shown in Eq. (9).
@@
Figure 4: Detailed operation of the proposed variable selection network.
As shown in Fig. 3(b) and Fig. 4, we propose a variable selection network to incorporate dynamic information. Auxiliary information and historical demand data used for urban EV charging demand prediction are inputted. The specific contribution of variables in each city area is determined by the network, i.e., variable selection network enables to emphasize variables that are critical to the prediction problem and ignore variables that are not independently at each node. The stacked input data $\bm{\xi}\in\mathbb{R}^{N_{j}\times\tau\times F}$ (including high-dimensional features $\bm{H}$, and other dynamics) are processed independently by two GRNs at feature dimension. A softmax function after one of the GRNs determines the variable selection weights, which are then multiplied with the embedded features. The calculation process of the variable selection network is shown in Eq. (11). Note that we do not operate on the node dimension and the timestep dimension, which means that attention can be allocated independently in spatio-temporal terms. In other words, the importance of dynamic features can be differentially determined.
@@
where $\bm{\xi}_{1}\in\mathbb{R}^{N_{j}\times\tau\times F}$ is the embedded features after the first step calculation, $\bm{v}_{1}\in\mathbb{R}^{N_{j}\times\tau\times F}$ is the variable selection weights, and $\bm{\xi}_{2}\in\mathbb{R}^{N_{j}\times\tau}$ denotes the output of variable selection network, which combines multiple features. $N_{j}$ denotes the number of city areas, $\tau$ denotes the number of time steps, and $F$ stands for the number of features.
Transformer utilizes self-attention mechanism to model sequences, breaking through the limitation that recurrent neural networks cannot be computed in parallel, and becoming a research hotspot for time series tasks. Considering the historical correlation and slight fluctuations in charging demand data, we adopt the framework of Transformer encoder and improve it by utilizing GRN and Gated Attention Network. Similarly to self-attention, we process the output of variable selection network (i.e., $\bm{\xi}_{2}$) at time step dimension independently with three GRNs to obtain the query, key, and value matrix $\bm{Q}$, $\bm{K}$, and $\bm{V}$.
The attention mechanism assigns weights to each input time step, and even irrelevant units are assigned small weights. This leads to two problems: 1) relevant units are assigned insufficient attention weights, and 2) noise units without contributions are assigned weights, leading to performance degradation. Considering that in long charging demand sequences not every interval is sufficiently helpful for prediction, we develop gated temporal attention, in which we perform a Scaled Dot-Product Attention operation, but Gumbel-Softmax is used as weight activation function instead of Softmax, as shown in Eq. (12) and Fig. 3(c).
@@ 4 Experiments
In this section, we first present the data description and the specific setup of the experiment. Then, we compare the proposed model with current state-of-the-art models to empirically illustrate the performance of CityEVCP. The model generalizability is also illustrated. Furthermore,we conduct ablation experiments to show that each module and each factor can provide a contribution to improving the prediction accuracy. Finally, the impact of hyperparameters is analyzed experimentally.
@@ 4.1 Data Description
The proposed model CityEVCP is validated on a benchmark dataset collected from June 19 to July 18, 2022 in Shenzhen, China. We divide the city into areas and update the occupancy rate of the counted charging piles within the area every 5 minutes. Consequently we can obtain 8640 timestamps for the month. The areas division and the number of charging piles included in each area are shown in this link [^1] https://github.com/IntelligentSystemsLab/ST-EVCDP. We divide the training set, validation set and test set by 6:1:3 in chronological order, i.e., 19 June to 6 July for training, 7 July to 9 July for validation and 10 July to 18 July for test. A total of 450,175 POI data are collected from Amap [^2] https://lbs.amap.com/. POIs include a total of 14 categories such as restaurant, shopping, scenic areas, etc. Temperature data is also collected from a weather station in Shenzhen, which has a sampling interval of 30 minutes. We use linear interpolation to enable a 5-minute interval in the temperature data, making it the same as the demand data.
@@ 4.2 Experimental Setup
The hyperparameters are chosen after a detailed search experiment and we provide hyperparameter analysis in Subsection 4.6. The final experimental setup we chose is as follows. First, the lookback window size $\tau$ is set to 12, i.e. we provide the model with 60 minutes of historical information. Second, in K-means clustering, we cluster the area into 10 classes. Third, the temperature coefficient $\varpi$ is set to 1.5 to not only provide sparse attention to temporal information, but also to avoid focusing attention on one-hot vector. Fourth, the number of encoder components $N_{E}$ is set to 2. Fifth, the maximum epoch numbers of training process is set to 2000, and model training will early stop if the loss of the validation set does not decrease within 50 epochs. Sixth, Mean Square Error (MSE) is used as the loss function for all comparison models. Seventh, we use Adam as an optimizer with a mini-batch size of 512. Seventh, we employ four metrics, i.e., Root Mean Squared Error (RMSE), R-Square Coefficient ($\rm{R}^{2}$), Relative Absolute Error (RAE), and Mean Absolute Error (MAE) to illustrate the performance of our proposed model. Finally, all experiments in this study are accomplished on a Windows personal computer containing a NVIDIA RTX 4000 GPU. We share the code for our proposed model CityEVCP in this link. [^3] https://github.com/kuanghx3/CityEVCP.
@@ 4.3 Comparison Experiment
In this subsection, the proposed model CityEVCP will be compared with representative models to illustrate the state-of-the-art performance. The models for comparison includes 3 sequential neural networks, 2 spatial neural networks and 5 spatio-temporal neural networks, which are described in detail as follows:
@@
  • TGCN [27]: A neural network-based traffic prediction method combining graph convolutional networks and gated recurrent units.
@@
  • BEGAN [30]: An air traffic demand prediction network combining long short-term memory and graph attention mechanisms.
  • FCSTGNN [51]: A fully-connected spatial-temporal graph neural network that captures integrated spatio-temporal dependencies in multivariate time series data by graph convolution operations.
  • HSTGCN [16]: A heterogeneous spatio-temporal graph convolutional network combining graph convolution and gated recurrent units for predicting EV charging demand at different spatio-temporal resolutions. In particular, it considers POI classification through separate fully connected networks.
@@
In Table 1, we demonstrate the performance of the above comparison models and CityEVCP in four prediction intervals (i.e., 15, 30, 45, and 60 minutes). The proposed model is the best in each prediction interval and on each metric. Compared to ten representative methods, our proposed method CityEVCP improves about 9.59% in RMSE, 10.90% in RAE, and 10.91% in MAE on average, and achieves the best fit in $\rm R^{2}$. Moreover, CityEVCP outperforms the ten comparison models by 18.50%, 10.41%, 8.31% and 7.97% on average in 15, 30, 45, and 60 minutes prediction tasks, respectively. CityEVCP performs well on relatively long-term prediction, which is a challenge in prediction problems, demonstrating the contribution of temporal modules (e.g., improved transformer encoder) and multi-source heterogeneous information to improve the accuracy of EV charging demand prediction. Specifically, we consider the same two time-series effects, price and temperature, in previously proposed multivariate time-series convolutional model, FCSTGNN, and in comparison, CityEVCP improves by 12.08% over FCSTGNN on average. Moreover, compared to the previous EV charging demand prediction method, HSTGCN, which considers points of interest in separate fully connected networks, our proposed model improves by 3.19%. This demonstrates that the effective introduction of external information such as the areas’ attributes and weather, as well as the improvement of the network structure can enhance the prediction accuracy of citywide EV charging demand.
To further illustrate the performance of the proposed model, we show the prediction results for two typical areas in Fig. 5. The prediction results of CityEVCP and HSTGCN and true area charging demand (i.e., occupancy) are shown. Although the proposed model CityEVCP does not perform perfectly at all times, it outperforms HSTGCN overall. The errors between two methods and the true label are also shown, in order to intuitively illustrate the performance of the model. Node 106 is a suburb located in Bao’an District, Shenzhen, with only six charging piles. Due to too few charging piles in the area, the demand for charging shows stability at some times of a day. CityEVCP is more willing to adjust than HSTGCN, even though the true labels do not change, which results in CityEVCP’s prediction accuracy being closer to the true values. This may be due to the ability of the proposed model to take into account dynamic auxiliary information, and reduce the inertia and lags. Node 183 is located in a residential area in Longgang District, Shenzhen and has 202 charging piles. CityEVCP has a smaller error than HSTGCN, illustrating its good performance in regular areas. CityEVCP copes well with small fluctuations in daily data because the model can enable sparse temporal attention and fully account for urban areas and dynamic influences. In contrast, HSTGCN tends to amplify data fluctuations and develop large prediction errors.
Figure 5: Lineplots of 60-min prediction results and errors for two example areas.
@@
In comparison models, firstly, GCN that only considers spatial adjacencies and lacks consideration of temporal factors is the worst. Second, complex models perform better than simple ones, because complex models can model nonlinear spatio-temporal relationships accurately. This is why current research favors delicate and complex networks, even though they may require more storage space and calculation. Thirdly, the consideration of influences on EV charging demand somehow brings about improvements in model performance. HGNN and HSTGCN, which incorporate area attribute information, have relatively outstanding performance in comparison models, illustrating that area attributes can help model spatial features from high dimensions and improve prediction accuracy by leveraging similar nodes. TFT and FCSTGNN, which combine dynamic influences, outperform models of equal volume in long-time prediction tasks (e.g.,45min and 60min), illustrating the important role of dynamic auxiliary information in enhancing the model’s forward-looking capability.
@@
We conduct additional experiments to show that our model does not just fit on a specific dataset. We randomly drop 20% of the 247 areas to create different geographic contexts in the experiments. To avoid randomness, the experiments are repeated five times and their results are averaged. On these new datasets, we selected one temporal model (i.e., LSTM) and three spatio-temporal models (i.e., BEGAN, FCSTGNN, and HSTGCN) as comparative models to demonstrate the performance of our proposed model. The results and average results for the 4 prediction intervals are showed in Table 2. In the additional experiments, the proposed model CityEVCP improves by 11.94% in RMSE, 12.95% in RAE and 13.50% in MAE on average compared to the comparison model, and achieves the best fit in $\rm R^{2}$.
@@
CityEVCP 2.69 4.08 5.05 5.83 4.41 92.81 84.50 77.09 69.63 81.01
Metrics (×$\times 10^{-2}$) RAE MAE